OpenStudy (anonymous):

Simplify: 3 square root 5 divided by 4 square root 5 A) 5 1/4 B) 5 1/12 C) 5 7/12 D) 5 4/3

OpenStudy (hossamhoussien):

\[\frac{ 3\sqrt{5} }{ 4\sqrt{5} }\] Is that what you meant?

OpenStudy (anonymous):

Yes

OpenStudy (albert0898):

Okay. We know that the roots are the same. When we divide and the terms are the same, we get 1. \[\frac{ 3\sqrt{5} }{ 4\sqrt{5} }\] We get 3/4 after dividing, which gets us .75

OpenStudy (anonymous):

Okay...